### v1

@@ 3 General relativity and the polar cap emission
Coherent radio emission from the polar caps of a neutron star enabled the discovery of pulsars. The pulsed emission profile originates from the cosmic lighthouse effect, which relies on the misalignment angle $\chi$ between the magnetic and rotation axis. Recent works have demonstrated the generation of pulsar radio emission from non-stationary pair plasma discharges (Philippov et al. 2020; Cruz et al. 2021). Rotation-induced vacuum gaps, regions of an unscreened parallel electric field, potentiate the acceleration of particles to very high energies and subsequent emission of gamma-ray photons via curvature radiation. These hard photons are absorbed in super-strong magnetic fields and create pairs, generating a QED cascade that populates the magnetosphere with a dense electron-positron plasma. The pair production bursts are generated at an angle to the local magnetic field, favouring the induction of electromagnetic modes (Cruz et al. 2021). However, these models rely on efficient particle acceleration along the polar cap field lines for QED processes to kick in. Beloborodov (2008) and Timokhin & Arons (2012) addressed the particle accelerator problem at the polar cap. These works identified the $\alpha\equiv j_{m}/j_{\mathrm{GJ}}$ parameter determinant in defining under which conditions particles accelerate up to relativistic energies and can produce secondary pairs. The motivation behind this parameter, which is the ratio between the electric current extracted from the polar cap and the local co-rotation current $j_{\mathrm{GJ}}=-\Omega_{*}\cdot B/2\pi$, is that when $10<\alpha<$ 1, the accelerator is inefficient because the extracted current is enough to sustain the twist of the magnetic field lines near the light cylinder, given by
$j_{m}=\frac{1}{4\pi}\left(\@vec{\nabla}\times\@vec{B}_{\mathrm{magnetosphere}}\right)_{\parallel}.$ (43)
When $\alpha>1$ or $\alpha<0$, charges extracted from the surface/atmosphere are insufficient to carry the magnetospheric required current. This current-starved state generates an electric field that accelerates particles. This flow becomes unstable to QED cascading, thus supplying the additional carriers. Under such conditions, the polar cap becomes an efficient accelerator, and pulsar radio emission is viable. However, three-dimensional particle-in-cell simulations of low obliquity rotators (i.e., $\chi\leq 40^{\circ}$) showed weak particle acceleration in the bulk of the polar region (Philippov et al. 2015b), confirming estimates using the $\alpha$ parameter (Bai & Spitkovsky 2010; Timokhin & Arons 2012). Therefore, low obliquity rotators could not generate pulsar radio emission, which is unsupported by observations.
@@
In the 3+1 formalism, we reformulate the $\alpha$ parameter analysis for the efficiency of the polar cap accelerator as the 4-norm of the current $j_{\mu}j^{\mu}$. Equivalently, we have an efficient accelerator for $j_{\mu}j^{\mu}>0$ (i.e., spacelike current) or inefficient accelerator for $j_{\mu}j^{\mu}<0$ (i.e., timelike current) (Belyaev & Parfrey 2016; Gralla et al. 2016; Huang et al. 2018). In this way, we can use the spacelike current as an observable for field lines that will sustain pair creation and are viable locations for the pulsar radio emission.
@@
$j_{\mu}j^{\mu}$ =${j}\cdot\@vec{j}-\left(\rho c\right)^{2}\approx{j^{\hat{r}}}^{2}-\left(\rho c\right)^{2}\approx\left(\rho_{\mathrm{GJ}}c\right)^{2}\left(\left(\frac{B^{\hat{\phi}}}{E^{\hat{\theta}}}\right)^{2}-1\right),$ (44)
@@
where we assumed the magnetospheric solution is close to the force-free regime where the charge density closely follows the co-rotation value $\rho\sim\rho_{\mathrm{GJ}}$ and the radial current approximates to the ratio between the azimuthal magnetic field and the polar component of the electric field, as obtained by Lyutikov (2011). Although simple, equation (44) is powerful enough to provide us with intuition on the interplay between magnetospheric plasma injection and general-relativistic effects. The first is made explicit through $B^{\hat{\phi}}$ that measures the torsion of the field lines, which is more efficient in the force-free regime where a strong poloidal current gets extracted from the polar cap. The second point relates to the frame-dragging effect (Beskin 1990; Muslimov & Tsygan 1992; Philippov et al. 2015a; Torres et al. 2023), which reduces the induced electric field due to the mismatch between the stellar and the spacetime rotation. In the force-free regime, the wind region possesses $B^{\hat{\phi}}\sim E^{\hat{\theta}}$, corresponding to the null electric current case. This case is not particularly interesting as this would mean that, in flat spacetime, the aligned rotator would not be able to emit in the radio band. However, the frame-dragging effect reduces $E^{\hat{\theta}}$ raising the ratio above unity and leading to the possibility of pulsar radio emission (Philippov et al. 2015a).
@@
To validate this intuitive picture drawn from equation (44), we simulated the neutron star magnetosphere with no pair production and injected plasma from the surface with parallel velocity as described in Sect. 2.7. This injection method has two advantages: (1) direct control over the extracted poloidal current limited by the cold magnetization parameter; (2) successfully reproduces the force-free magnetospheric solution. Simulations ran in the half-domain setup for $R_{s}/R_{*}\in\left[0.0,0.6\right]$ and $\sigma_{*}\in\left[1000,2000\right]$, presented in Fig. 3. As the maximum meridional extension predicted for the spacelike current region (Belyaev & Parfrey 2016; Gralla et al. 2016), $\theta_{\mathrm{SL}}$, is a function of the stellar compactness, the meridional resolution changed accordingly for each simulation to resolve this region with at least 20 numerical cells. This angle also changes with the stellar angular velocity, fixed at $\Omega_{*}=0.1\leavevmode\nobreak\ [\mathrm{c\leavevmode\nobreak\ rad/R_{*}}]$ for this set of simulations. Table 1 contains the detailed list of parameters used.
@@
The first striking feature in Fig. 3 is the inexistence of the spacelike region for some non-zero compactness simulations. The explanation lies in the poloidal current and, consequently, the degree of field line torsion. When $\sigma_{*}=1000$, the extracted poloidal current is very close to the expected for the force-free split-monopole case, thus ensuring $B^{\hat{\phi}}\sim E^{\hat{\theta}}$. In this way, any general-relativistic reduction of the electric field results in a spacelike current, following equation (44). Lower plasma supply cases represented by higher magnetizations yield a decrease in the poloidal current extracted from the polar cap. The $B^{\hat{\phi}}$ reduction no longer guarantees the existence of the spacelike region at any compactness, thus explaining the appearance of cut-off conditions as shown in Fig. 4.
@@
From the observational point of view, this shows the robustness of the pulsar radio emission mechanism at any inclination. We have just shown that the spacelike region appears at any compactness for force-free magnetospheric solutions, even for the least favourable inclination configuration (the aligned rotator). Nevertheless, the dependence of the radio beam generation mechanism on the overall magnetospheric solution could explain why some neutron stars are radio-quiet or may present intermittent behaviour. Weak pulsars or pulsars near the death line, where pair production is no longer very efficient, could be strong candidates for this scenario.
@@
Run $\mu_{*}\leavevmode\nobreak\ \left[\mathrm{m_{e}c^{2}R_{*}^{2}/e}\right]$ $\Omega_{*}\leavevmode\nobreak\ \left[\mathrm{c\leavevmode\nobreak\ rad/R_{*}}\right]$ $R_{s}\leavevmode\nobreak\ \left[\mathrm{R_{*}}\right]$ $\sigma_{*}$ Injection $n_{\mathrm{inj}}$ $v_{\parallel}\leavevmode\nobreak\ \left[\mathrm{c}\right]$ Resolution Spacelike region
@@
Table 1: Simulation parameters used for the first numerical experiment (Sect. 3), shown in Figs. 3 and 4.
@@
Figure 4: Maximum poloidal extension of the spacelike region ($\theta_{\mathrm{SL}}$) as function of the stellar compactness $R_{s}/R_{*}$. The grey line corresponds to the force-free estimate in Gralla et al. (2016). The circles correspond to the maximum and minimum measured angles extracted from numerical simulations presented in Fig. 3. This figure captures the transition from inactive to active polar caps with increasing compactness, e.g. for the $\sigma_{*}=1500\leavevmode\nobreak\ (2000)$ case, the polar cap activates for $R_{s}/R_{*}\gtrsim 0.4\leavevmode\nobreak\ (0.5)$.
@@ 4 Transition from weak to energetic pulsars
In the previous section, we validated the intuitive picture given by equation (44). However, we adopted a controlled injection methodology that drove the system to the desired solution. In this section, we improve the magnetospheric model by capturing the gap dynamics through heuristic ”on-the-spot” pair production (Cruz et al. 2022), i.e. in the limit of vanishing photon mean free path. Under these circumstances, solutions become sensitive to the upstream gap plasma supply from the neutron star surface. The charges lifted from the surface form an atmosphere maintained as described in Sect. 2.7, feeding self-consistently the magnetosphere with cold plasma. We follow the PSG non-stationary gap model by sustaining a minimum atmospheric plasma density at the base $n_{0}\equiv n_{+}+n_{-}\sim 2n_{\mathrm{GJ}}$. A weak parallel component of the electric field persists at the base of some magnetic field lines, providing an initial kick to those particles extracted from the atmosphere.
Figure 5 shows the quasi-stationary state of a millisecond neutron star with compactness $R_{s}/R_{*}=0.3$ and angular velocity $\Omega_{*}=0.2\leavevmode\nobreak\ \left[\mathrm{c\leavevmode\nobreak\ rad/R_{*}}\right]$ for different pair production efficiencies $\eta$ given by:
@@
where $\gamma_{\mathrm{max}}\sim B_{*}\Omega_{*}^{2}$ is the estimated maximum achievable particle Lorentz factor. The magnetic field amplitude controls the pair production efficiency by keeping constant the threshold and secondary pair energy at $\gamma_{\mathrm{thr}}=25\leavevmode\nobreak\ [m_{e}c^{2}]$ and $\gamma_{\mathrm{pair}}=8\leavevmode\nobreak\ [m_{e}c^{2}]$. Also, we restrict pair production to be active only up to three stellar radii (i.e., $R_{\mathrm{PP}}=3R_{*}$). Table 2 summarizes the simulation parameters used in this numerical experiment. The low pair production efficiency case, with $\eta\sim 38$, highlights the typical solution of a weak pulsar with a persistent outer gap (Shinya & Shinpei 2012; Gruzinov 2012, 2013, 2015; Chen et al. 2020). This gap is a consequence of the inability of the surface and the pair production in the volumetric return current to supply plasma to the region behind the null surface, i.e. the region where the poloidal magnetic field lines curve towards the equator. As the efficiency increases, the amplitude of the gap’s parallel electric field reduces until it becomes almost zero, when the magnetosphere transitions to a force-free solution. The volumetric return current that occupied a portion of the polar cap for low-efficiency values now occupies lower latitudes, increasing the effective polar cap area up to the maximum poloidal extension angle $\theta_{\mathrm{PC}}$. Consequently, the magnetosphere can increase the electric current driven through the open flux tube, leading to a jump in the Poynting flux luminosity. In this model, and for simplicity, we kept the pair production efficiency constant in the magnetosphere. However, different QED processes play a role in distinct magnetospheric regions (Chen & Beloborodov 2014), which may lead to differential pair production efficiencies (deferred for future work). Magnetospheres that continuously oscillate between weak and quasi-force-free states, depending on the morphology of the outer gap, can explain the observed phenomenon of pulsar luminosity intermittency (similar to Li et al. (2012)).
@@
Several crucial changes accompany the transition from weak to force-free solutions: 1) the Y-point moves closer to the light-cylinder radius due to the more efficient screening of the equatorial superrotation region (Shinya & Shinpei 2012; Hu & Beloborodov 2022); 2) the stronger poloidal current driven through the magnetosphere induces a higher torsion of the magnetic field lines in the toroidal direction. The latter is vital for the activation of the polar cap. At fixed compactness (in this case $R_{s}/R_{*}=0.3$), Fig. 4 shows that the polar cap may present a spacelike volumetric region only if the magnetosphere drives a stronger current, characteristic of higher plasma supply scenarios. With heuristic pair production, the efficiency controls the plasma supplied from the polar cap gap to the magnetosphere, thus having the same effect as the surface plasma supply from the previous section. In Fig. 5, we observe the appearance of the spacelike volume in the polar cap for efficiencies $\eta\gtrsim 155$. However, many processes affect this minimum efficiency value $\eta_{\mathrm{min}}$: 1) the frame-dragging effect for higher compactness neutron stars reduces $\eta_{\mathrm{min}}$ due to the reduction of the poloidal electric field; 2) evident interplay between the outer and inner polar cap gap, more efficient pair production at the Y-point (e.g., including photon-photon processes) pulls the return current away from the polar cap thus increasing the field line torsion of the magnetosphere, which reduces $\eta_{\mathrm{min}}$; 3) a denser neutron star atmosphere also facilitates the launch of a stronger poloidal current, in a similar way as in point 2. In particular, point 2 suggests a correlation between the radio beam and the existence of the outer gap. Intermittent pulsars may also present intermittent radio beams, potential candidates for explaining sub-pulse nulling.
@@
Solutions with very low pair production efficiencies are sustained solely by pair plasma generated at the return current region and progressively tend to the ”dead pulsar” solution (electrosphere), corresponding to the no pair production case Cruz et al. (2023). Solutions with even higher pair production efficiencies than the ones presented in this study tend to the force-free regime and possess an active polar cap, i.e. in the accelerator regime. In this sense, we retrieve the same conclusions as in the previous section: low pair production efficiencies (i.e., low plasma supply) require higher stellar compactnesses for the activation of the pulsar mechanism; higher pair production efficiencies (i.e., high plasma supply) relax the compactness restriction. Nevertheless, for low-obliquity rotators, compactness is crucial in the generation phase of the radio beam along the magnetic axis.
@@
Figure 5: Panels showing the positronic charge density ($\rho_{+}$), 4-norm of the current ($j_{\mu}j^{\mu}$), FIDO-measured radial current ($J^{\hat{r}}$), and parallel component of the electric field ($E_{\parallel}$) for increasing pair production efficiency ($\eta$). Field components are multiplied by the radial distance squared ($r^{2}$) to highlight magnetospheric features further away from the star. Magnetic field lines are in grey and the magnetic field lines that start at the maximum spacelike angle ($\theta_{\mathrm{SL}}$) and polar cap extension angle ($\theta_{\mathrm{PC}}$) are shown in red. This figure captures the transition from a weak to force-free solution, obtained by increasing the stellar magnetic moment and, hence, the pair production efficiency. Consequently, the Y-point approaches the $R_{\mathrm{LC}}$, the polar cap extends to $\theta_{\mathrm{PC}}$, the parallel electric field component almost vanishes everywhere, and the spacelike current and positrons from pair production appear in the polar cap.

### v2

@@ 3 General relativity and the polar cap emission
Coherent radio emission from the polar caps of a neutron star enabled the discovery of pulsars. The pulsed emission profile originates from the cosmic lighthouse effect, which relies on the misalignment angle $\chi$ between the magnetic and rotation axis. Recent works have demonstrated the generation of pulsar radio emission from non-stationary pair plasma discharges (Philippov et al., 2020; Cruz et al., 2021). Rotation-induced vacuum gaps, regions of an unscreened parallel electric field, potentiate the acceleration of particles to very high energies and subsequent emission of gamma-ray photons via curvature radiation. These hard photons are absorbed in super-strong magnetic fields and create pairs, generating a QED cascade that populates the magnetosphere with a dense electron-positron plasma. The pair production bursts are generated at an angle to the local magnetic field, favouring the induction of electromagnetic modes (Cruz et al., 2021). However, these models rely on efficient particle acceleration along the polar cap field lines for QED processes to kick in. Beloborodov (2008) and Timokhin and Arons (2012) addressed the particle accelerator problem at the polar cap. These works identified the $\alpha\equiv j_{m}/j_{\mathrm{GJ}}$ parameter determinant in defining under which conditions particles accelerate up to relativistic energies and can produce secondary pairs. The motivation behind this parameter, which is the ratio between the electric current extracted from the polar cap and the local co-rotation current $j_{\mathrm{GJ}}=-\Omega_{*}\cdot B/2\pi$, is that when $10<\alpha<$ 1, the accelerator is inefficient because the extracted current is enough to sustain the twist of the magnetic field lines near the light cylinder, given by
$j_{m}=\frac{1}{4\pi}\left(\boldsymbol{\nabla}\times\boldsymbol{B}_{\mathrm{magnetosphere}}\right)_{\parallel}.$ (43)
When $\alpha>1$ or $\alpha<0$, charges extracted from the surface/atmosphere are insufficient to carry the magnetospheric required current. This current-starved state generates an electric field that accelerates particles. This flow becomes unstable to QED cascading, thus supplying the additional carriers. Under such conditions, the polar cap becomes an efficient accelerator, and pulsar radio emission is viable. However, three-dimensional particle-in-cell simulations of low obliquity rotators (i.e., $\chi\leq 40^{\circ}$) showed weak particle acceleration in the bulk of the polar region (Philippov et al., 2015b), confirming estimates using the $\alpha$ parameter (Bai and Spitkovsky, 2010; Timokhin and Arons, 2012). Therefore, low obliquity rotators could not generate pulsar radio emission, which is unsupported by observations.
@@
In the 3+1 formalism, we reformulate the $\alpha$ parameter analysis for the efficiency of the polar cap accelerator as the 4-norm of the current $j_{\mu}j^{\mu}$. Equivalently, we have an efficient accelerator for $j_{\mu}j^{\mu}>0$ (i.e., spacelike current) or inefficient accelerator for $j_{\mu}j^{\mu}<0$ (i.e., timelike current) (Belyaev and Parfrey, 2016; Gralla et al., 2016; Huang et al., 2018). In this way, we can use the spacelike current as an observable for field lines that will sustain pair creation and are viable locations for the pulsar radio emission. Therefore, characterizing the spacelike current region may lead to the direct characterization of the emitted polar radio beam, bridging the gap between numerical models and observations.
@@
$j_{\mu}j^{\mu}$ $=\boldsymbol{j}\cdot\boldsymbol{j}-\left(\rho c\right)^{2}\approx{j^{\hat{r}}}^{2}-\left(\rho c\right)^{2}\approx\left(\rho_{\mathrm{GJ}}c\right)^{2}\left(\left(\frac{B^{\hat{\phi}}}{E^{\hat{\theta}}}\right)^{2}-1\right),$ (44)
@@
where we assumed the magnetospheric solution is close to the force-free regime where the charge density closely follows the co-rotation value $\rho\sim\rho_{\mathrm{GJ}}$ and the radial current approximates to the ratio between the azimuthal magnetic field and the polar component of the electric field, as obtained by Lyutikov (2011). Although simple, equation (44) is powerful enough to provide us with intuition on the interplay between magnetospheric plasma injection and general-relativistic effects. The first is made explicit through $B^{\hat{\phi}}$ that measures the torsion of the field lines, which is more efficient in the force-free regime where a strong poloidal current gets extracted from the polar cap. The second point relates to the frame-dragging effect (Beskin, 1990; Muslimov and Tsygan, 1992; Philippov et al., 2015a; Torres et al., 2023), which reduces the induced electric field due to the mismatch between the stellar and the spacetime rotation. In the force-free regime, the wind region possesses $B^{\hat{\phi}}\sim E^{\hat{\theta}}$, corresponding to the null electric current case. This case is not particularly interesting as this would mean that, in flat spacetime, the aligned rotator would not be able to emit in the radio band. However, the frame-dragging effect reduces $E^{\hat{\theta}}$ raising the ratio above unity and leading to the possibility of pulsar radio emission (Philippov et al., 2015a).
@@
To validate this intuitive picture drawn from equation (44), we simulated the neutron star magnetosphere with no pair production and injected plasma from the surface with parallel velocity as described in Sect. 2.7. This injection method has two advantages: (1) direct control over the extracted poloidal current limited by the cold magnetization parameter; (2) successfully reproduces the force-free magnetospheric solution. Simulations ran in the half-domain setup for $R_{s}/R_{*}\in\left[0.0,0.6\right]$ and $\sigma_{*}\in\left[1000,2000\right]$, presented in Fig. 4. As the maximum meridional extension predicted by force-free models for the spacelike current region (Belyaev and Parfrey, 2016; Gralla et al., 2016), $\theta_{\mathrm{SL}}$, is a function of the stellar compactness, the meridional resolution changed accordingly for each simulation to resolve this region with at least 20 numerical cells. This angle also changes with the stellar angular velocity, fixed at $\Omega_{*}=0.1~{}[\mathrm{c~{}rad/R_{*}}]$ for this set of simulations. Table 1 contains the detailed list of parameters used. It is important to reiterate that our study takes into account all general-relativistic effects, not just the frame-dragging, unlike previous works on neutron stars (e.g., Philippov and Spitkovsky (2018); Chen et al. (2020); Bransgrove et al. (2022)). In fact, as is shown in Fig. 3, this approximation leads to an overestimation of both the polar cap and spacelike angles of $\sim 21.8\%$ for $R_{s}/R_{*}=0.5$ (typical stellar compactness selected), following the general-relativistic force-free prediction given in Gralla et al. (2016). This overestimation leads to macroscopic magnetospheric changes, including an increased effective open flux tube area, spin-down luminosity, and wider generated radio beam. This point is particularly important when simulating magnetospheric environments of compact neutron stars approaching the Tolman–Oppenheimer–Volkoff mass limit (Kalogera and Baym, 1996; Romani et al., 2022) or long-lived hypermassive neutron stars (Falcke and Rezzolla, 2014; Chirenti et al., 2019, 2023; Selvi et al., 2024).
Figure 3: Force-free estimates of the maximum polar cap, $\theta_{\mathrm{PC}}$, and spacelike, $\theta_{\mathrm{SL}}$, poloidal extension at the neutron star surface for varying stellar compactness values, $R_{s}/R_{*}$. Curves are obtained using the model developed in Gralla et al. (2016) for a general-relativistic description of force-free electrodynamics. The complete description is presented in blue. Estimates neglecting the frame-dragging contribution and the lapse function are presented in green and orange, respectively. We conclude that not considering the lapse function leads to an overestimate of the open flux tube and effective particle acceleration volume, thus a macroscopic inaccuracy.
@@
The first striking feature in Fig. 4 is the inexistence of the spacelike region for some non-zero compactness simulations. The explanation lies in the poloidal current and, consequently, the degree of field line torsion. When $\sigma_{*}=1000$, the extracted poloidal current is very close to the expected for the force-free split-monopole case, thus ensuring $B^{\hat{\phi}}\sim E^{\hat{\theta}}$. In this way, any general-relativistic reduction of the electric field results in a spacelike current, following equation (44). Lower plasma supply cases represented by higher magnetizations yield a decrease in the poloidal current extracted from the polar cap. The $B^{\hat{\phi}}$ reduction no longer guarantees the existence of the spacelike region at any compactness, thus explaining the appearance of cut-off conditions as shown in Fig. 5. Figure 5 highlights the first direct comparison between global particle-in-cell simulations and force-free model predictions for the meridional extension of the spacelike current volume. Results show a remarkable agreement in the force-free limit but start to deviate for more charge-starved magnetospheric solutions. Nevertheless, the force-free estimates limit the predicted spacelike extension from above, thus it is still a good estimate for the maximum generated radio beam width.
@@
From the observational point of view, this shows the robustness of the pulsar radio emission mechanism at any inclination. We have just shown that the spacelike region appears at any compactness for force-free magnetospheric solutions, even for the least favourable inclination configuration (the aligned rotator). Nevertheless, the dependence of the radio beam generation mechanism on the overall magnetospheric solution could explain why some neutron stars are radio-quiet or may present intermittent behaviour. Weak pulsars or pulsars near the death line, where pair production is no longer very efficient, could be strong candidates for this scenario. Also, our model predicts narrower radio beam widths for older low-obliquity pulsars, whose magnetospheres are more charge-starved. In some cases, this width reduction can reach half the force-free value, which may be measurable through observations. In particular, observations of millisecond pulsars report a systematic tendency for smaller inferred luminosities and narrower emission beams (Kramer et al., 1998), which support our findings. Thus, the compactness of millisecond pulsar magnetospheres and deviations from force-free solutions may explain the reduction of the angular radius of the generated radio beams.
@@
Run $\mu_{*}~{}\left[\mathrm{m_{e}c^{2}R_{*}^{2}/e}\right]$ $\Omega_{*}~{}\left[\mathrm{c~{}rad/R_{*}}\right]$ $R_{s}~{}\left[\mathrm{R_{*}}\right]$ $\sigma_{*}$ Injection $n_{\mathrm{inj}}$ $v_{\parallel}~{}\left[\mathrm{c}\right]$ Resolution Spacelike region
@@
Table 1: Simulation parameters used for the first numerical experiment (Sect. 3), shown in Figs. 4 and 5. These simulations take on average 200k CPU core hours.
@@
Figure 5: Maximum poloidal extension of the spacelike region ($\theta_{\mathrm{SL}}$) as function of the stellar compactness $R_{s}/R_{*}$. The grey line corresponds to the force-free estimate in Gralla et al. (2016). The circles correspond to the maximum and minimum measured angles extracted from numerical simulations presented in Fig. 4. This figure captures the transition from inactive to active polar caps with increasing compactness, e.g. for the $\sigma_{*}=1500~{}(2000)$ case, the polar cap activates for $R_{s}/R_{*}\gtrsim 0.4~{}(0.5)$.
@@ 4 Transition from weak to energetic pulsars
In the previous section, we validated the intuitive picture given by equation (44). However, we adopted a controlled injection methodology that drove the system to the desired solution. In this section, we improve the magnetospheric model by capturing the gap dynamics through heuristic “on-the-spot” pair production (Cruz et al., 2022), i.e. in the limit of vanishing photon mean free path. Under these circumstances, solutions become sensitive to the upstream gap plasma supply from the neutron star surface. The charges lifted from the surface form an atmosphere maintained as described in Sect. 2.7, feeding self-consistently the magnetosphere with cold plasma. We follow the PSG non-stationary gap model by sustaining a minimum atmospheric plasma density at the base $n_{0}\equiv n_{+}+n_{-}\sim 2n_{\mathrm{GJ}}$. A weak parallel component of the electric field persists at the base of some magnetic field lines, providing an initial kick to those particles extracted from the atmosphere.
Figure 6 shows the quasi-stationary state of a millisecond neutron star with compactness $R_{s}/R_{*}=0.3$ (e.g., Bhattacharyya et al. (2017)) and angular velocity $\Omega_{*}=0.2~{}\left[\mathrm{c~{}rad/R_{*}}\right]$ for different pair production efficiencies $\eta$ given by:
@@
where $\gamma_{\mathrm{max}}\sim B_{*}\Omega_{*}^{2}$ is the estimated maximum achievable particle Lorentz factor. The magnetic field amplitude controls the pair production efficiency by keeping constant the threshold and secondary pair energy at $\gamma_{\mathrm{thr}}=25~{}[m_{e}c^{2}]$ and $\gamma_{\mathrm{pair}}=8~{}[m_{e}c^{2}]$. Also, we restrict pair production to be active only up to three stellar radii (i.e., $R_{\mathrm{PP}}=3R_{*}$). Table 2 summarizes the simulation parameters used in this numerical experiment. The low pair production efficiency case, with $\eta\sim 38$, highlights the typical solution of a weak pulsar with a persistent outer gap (Shinya and Shinpei, 2012; Gruzinov, 2012, 2013, 2015; Chen et al., 2020). This gap is a consequence of the inability of the surface and the pair production in the volumetric return current to supply plasma to the region behind the null surface, i.e. the region where the poloidal magnetic field lines curve towards the equator. As the efficiency increases, the amplitude of the gap’s parallel electric field reduces until it becomes almost zero, when the magnetosphere transitions to a force-free solution. The volumetric return current that occupied a portion of the polar cap for low-efficiency values now occupies lower latitudes, increasing the effective polar cap area up to the maximum poloidal extension angle $\theta_{\mathrm{PC}}$, as predicted by force-free models (Gralla et al., 2016; Belyaev and Parfrey, 2016). Consequently, the magnetosphere can increase the electric current driven through the open flux tube, leading to a jump in the Poynting flux luminosity. In this model, and for simplicity, we kept the pair production efficiency constant in the magnetosphere. However, different QED processes play a role in distinct magnetospheric regions (Chen and Beloborodov, 2014), which may lead to differential pair production efficiencies (deferred for future work). Magnetospheres that continuously oscillate between weak and quasi-force-free states, depending on the morphology of the outer gap, can explain the observed phenomenon of pulsar luminosity intermittency (similar to Li et al. (2012)).
@@
Several crucial changes accompany the transition from weak to force-free solutions: 1) the Y-point moves closer to the light-cylinder radius due to the more efficient screening of the equatorial superrotation region (Shinya and Shinpei, 2012; Hu and Beloborodov, 2022); 2) the stronger poloidal current driven through the magnetosphere induces a higher torsion of the magnetic field lines in the toroidal direction. The latter is vital for the activation of the polar cap. At fixed compactness (in this case $R_{s}/R_{*}=0.3$), Fig. 5 shows that the polar cap may present a spacelike volumetric region only if the magnetosphere drives a stronger current, characteristic of higher plasma supply scenarios. With heuristic pair production, the efficiency controls the plasma supplied from the polar cap gap to the magnetosphere, thus having the same effect as the surface plasma supply from the previous section. In Fig. 6, we observe the appearance of the spacelike volume in the polar cap for efficiencies $\eta\gtrsim 155$. However, many processes affect this minimum efficiency value $\eta_{\mathrm{min}}$: 1) the frame-dragging effect for higher compactness neutron stars reduces $\eta_{\mathrm{min}}$ due to the reduction of the poloidal electric field; 2) evident interplay between the outer and inner polar cap gap, more efficient pair production at the Y-point (e.g., including photon-photon processes) pulls the return current away from the polar cap thus increasing the field line torsion of the magnetosphere, which reduces $\eta_{\mathrm{min}}$; 3) a denser neutron star atmosphere also facilitates the launch of a stronger poloidal current, in a similar way as in point 2. In particular, point 2 suggests a correlation between the radio beam and the existence of the outer gap. Intermittent pulsars may also present intermittent radio beams, potential candidates for explaining pulse nullings. We have also performed simulations where we inject an atmospheric plasma composed of electrons and ions, where the ions have the same mass and charge as positrons but are not allowed to pair produce (Philippov and Spitkovsky, 2018). As expected, the injection of ions leads to a lower pair plasma supply to the outer magnetosphere of weak pulsars, thus increasing the minimum efficiency value $\eta_{\mathrm{min}}$. Apart from a shift of the magnetospheric solution towards charge-starvation, we do not observe significant changes in the outer gap dynamics or magnetospheric structure. However, this point may change if ions possess realistic $q/m$, which we differ to a future study. Also, the impact of ion injection on the magnetosphere is more significant in charge-starved pulsar solutions due to the high degree of charge separation, thus this effect is diluted for denser solutions approaching the force-free regime.
@@
Solutions with very low pair production efficiencies are sustained solely by pair plasma generated at the return current region and progressively tend to the “dead pulsar” solution (electrosphere), corresponding to the no pair production case (Cruz et al., 2023). Solutions with even higher pair production efficiencies than the ones presented in this study tend to the force-free regime and possess an active polar cap, i.e. in the accelerator regime. In this sense, we retrieve the same conclusions as in the previous section: low pair production efficiencies (i.e., low plasma supply) require higher stellar compactnesses for the activation of the pulsar mechanism; higher pair production efficiencies (i.e., high plasma supply) relax the compactness restriction. Nevertheless, for low-obliquity rotators, compactness is crucial in the generation phase of the radio beam along the magnetic axis. Also, as observed in the previous section, the width of the radio beam deviates from the force-free estimate depending on the charge-starvation state. Thus, for older and weaker low-obliquity pulsars, the radio beam width should be reduced.
Another important feature on the structure of the magnetosphere is the location of the Y-point, which for all simulations presented is located at a significant fraction of the light-cylinder radius (i.e. $r_{\mathrm{Y}}/R_{\mathrm{LC}}\sim 0.6-0.8$), consistent with previous kinetic models of neutron star magnetospheres (Chen et al., 2020; Guépin et al., 2020; Cruz et al., 2023; Hakobyan et al., 2023). The spatial restriction imposed on the heuristic pair production ($R_{\mathrm{PP}}=3R_{*}$) limits the supply of plasma to the outer magnetosphere, in particular, to regions close to the Y-point. This weak supply of plasma beyond $R_{\mathrm{LC}}$ explains the formation of the persistent outer gap, volumetric return current and wide equatorial current sheet. This is in contrast to the typical ideal force-free magnetosphere with abundant plasma supply, clear and localized Y-shaped return current and thin equatorial current sheet. Wider current sheets dissipate energy at a much slower rate, thus we expect our novel reduced-domain model (i.e., imposing equatorial symmetry) to be accurate in modelling weak pulsar magnetospheres in the aligned rotator configuration. We have also verified the convergence of the presented results by doubling the number of particles per cell injected every timestep.
Particularly interesting is the systematic formation of core component-associated beamlets at lower pair production efficiencies (see positronic charge density for $\eta=38-77$ simulations in Fig. 6). These beamlets are generated by the cyclic opening of a fragile gap near the magnetic pole of the star. As pair production is prohibited close to the axis due to the small curvature of the magnetic field lines, some electrons from the magnetosphere are reversed back to the star, thus intermittently shorting out the accelerating gap from above (Lyubarsky, 2009). This cyclic behaviour allows the magnetosphere to drive the required quasi-steady-state polar current that supports the magnetic field line torsion in the outer magnetosphere. We conjecture that these beamlets are responsible for the narrow dwarf pulses observed during ordinary pulse nullings of PSR B2111+46 in Chen et al. (2023), referred to as “particle raindrops” by the authors.
@@
Figure 6: Panels showing the positronic charge density ($\rho_{+}$), 4-norm of the current ($j_{\mu}j^{\mu}$), FIDO-measured radial current ($J^{\hat{r}}$), and parallel component of the electric field ($E_{\parallel}$) for increasing pair production efficiency ($\eta$). Field components are multiplied by the radial distance squared ($r^{2}$) to highlight magnetospheric features further away from the star. Magnetic field lines are in grey and the magnetic field lines that start at the maximum spacelike angle ($\theta_{\mathrm{SL}}$) and polar cap extension angle ($\theta_{\mathrm{PC}}$) are shown in red. This figure captures the transition from a weak to force-free solution, obtained by increasing the stellar magnetic moment and, hence, the pair production efficiency. Consequently, the Y-point (at $r/R_{*}\sim r_{\mathrm{Y}}$) approaches the $R_{\mathrm{LC}}$, the polar cap extends to $\theta_{\mathrm{PC}}$, the parallel electric field component almost vanishes everywhere, and the spacelike current and positrons from pair production appear in the polar cap (see $j_{\mu}j^{\mu}$ zoom insets). The location of the low-multiplicity beamlets is highlighted in the positronic charge density insets close to the poloidal axes.